How much electrical energy is required to generate 1.000L of H, at 25°C and 1.00 atm pressure by the electrolysis of water?


1
Expert's answer
2021-01-19T04:36:02-0500

Pv = nRT .

1 × 1 = n × 0.0821 × 298 .

n = 1 / 0.0821 × 298 .

n = 0.0408 mol. Mol of H2 gas .


For

2 H+ + 2e ----------- H2.




For 1mol of H2 2mol of e required .




So , mol of electron required = 2n = 2 × 0.04

= 0.08 mol of electron .


Charge = ( 0.08 × 96500 ) Coulomb.

= 7720 C .